I have 20 of these questions to do, so once I get help with this first one, I beleive I shall be able to get the other 19 out. Thank you so much if you can help me. (My email is sasha_mahabir@hotmail.com)
HERE IS THE QUESTION:
1.) Write an algebraic function for each of the following. Use 'y' as the dependant variable and 'x' as the independant varaiable. Then evaluate the functions for the input values 9, 16.37, 0, and -10.
a.) The output is the opposite of the input.
b.) The output is the product of five and the sum of the input and six.
c.) The input is the sum of five and the product of the input and six.
d.) The output is always 100.
e.) The output is the square of the product of 3 and the input.
f.) The output is the product of three and the square of the input.
Answer by stanbon(75887) 
 1.) Write an algebraic function for each of the following. Use 'y' as the dependant variable and 'x' as the independant varaiable. Then evaluate the functions for the input values 9, 16.37, 0, and -10.
a.) The output is the opposite of the input.
y=-x
If x=9, y=-9
If x=16, y=-16
etc.
b.) The output is the product of five and the sum of the input and six.
y=5(x+6)
i x=9, y=5(9+6), y=5*15, y=75
etc.
c.) The output is the sum of five and the product of the input and six.
y=5+6x
If x=0 y=5+6*9, y=5+54, y=59
etc.
d.) The output is always 100.
y=100
If x=9, y=100
etc.
e.) The output is the square of the product of 3 and the input.
y=(3x)^2
If x=9, y=(3*9)^2, y= 27^2, y= 729
etc.
f.) The output is the product of three and the square of the input.
y=3x^2
If x=9,y=3*9^2, y=3*81, y=243
etc.
==============================
Cheers,
Stan H.
|
|
|
| |